# Quantum Dynamic Programming

### v1

@@ II Locally accurate implementations
Consider a quantum recursion defined by Eq. (7) with the root state $\sigma_{0}$. The exact implementation of memory-calls in recursion steps would lead to a sequence of states $\rho_{0},\rho_{1},\ldots$. We need to consider the situation where each recursion step consists of approximations of memory-calls.
@@
$\frac{1}{2}\left\lVert\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{k},M)}_{\text{QDP}}({\sigma_{k}})-\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\sigma_{k})}(\sigma_{k})\right\rVert_{1}=\frac{1}{2}\left\lVert\sigma_{k+1}-\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\sigma_{k})}(\sigma_{k})\right\rVert_{1}\leq\epsilon.$ (19)
@@
For clarity we assume the recursion step consists of a single memory-call implemented by $M$ memory-usage queries. We then have that the ideal implementation is $\rho_{1}=\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho_{0})}(\rho_{0})=\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\sigma_{0})}(\sigma_{0})$ and the practical implementation $\sigma_{1}=\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{0},M)}_{\text{QDP}}(\sigma_{0})$, such that
$\Xi_{1}$ :-$\coloneq\frac{1}{2}\|\sigma_{1}-\rho_{1}\|_{1}=\frac{1}{2}\left\|\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{0},M)}_{\text{QDP}}(\sigma_{0})-\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\sigma_{0})}(\sigma_{0})\right\|_{1}=\mathcal{O}(1/M).:$ (20)
@@
$\frac{1}{2}\left\|\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{0},M)}_{\text{QDP}}-\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\sigma_{0})}\right\|_{\operatorname{Tr}}=\mathcal{O}(1/M)\ .$ (21)
The QDP channel ^$\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{1},M)}_{\text{QDP}}$ consists of $M$ memory-usage queries. For each such memory-usage query defined by
^s$\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{1})}_{s}(\omega)=\operatorname{Tr}_{1}\left[e^{-i\hat{N}s}\left(\sigma_{1}\otimes\omega\right)e^{i\hat{N}s}\right]\ ,$ (22)
@@
^s$\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{1})}_{s}(\omega)=\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho_{1})}_{s}(\omega)+\operatorname{Tr}_{1}\left[e^{-i\hat{N}s}\left((\sigma_{1}-\rho_{1})\otimes\omega\right)e^{i\hat{N}s}\right]\ ,$ (23)
@@
$\frac{1}{2}\left\|\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{1})}_{s}-\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho_{1})}_{s}\right\|_{\operatorname{Tr}}\leq\frac{1}{2}\|\sigma_{1}-\rho_{1}\|_{1}=\Xi_{1}\ .$ (24)
Here we see how the preparation error $\Xi_{1}$ of $\sigma_{1}$ compared to $\rho_{1}$ influences the distance of the recursion step maps that define $\sigma_{2}=\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{1},M)}_{\text{QDP}}(\sigma_{1})$ and $\rho_{2}=\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho_{1},M)}_{\text{QDP}}(\rho_{1})$ so that using telescoping and the triangle inequality
$\frac{1}{2}\left\|\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{1},M)}_{\text{QDP}}-\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho_{1},M)}_{\text{QDP}}\right\|_{\operatorname{Tr}}\leq\frac{M}{2}\left\|\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{1})}_{1/M}-\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho_{1})}_{1/M}\right\|_{\operatorname{Tr}}=\mathcal{O}(M\Xi_{1})\ .$ (25)
@@
$\Xi_{2}$ $=\frac{1}{2}\left\|\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{1},M)}_{\text{QDP}}(\sigma_{1})-\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho_{1})}(\rho_{1})\right\|_{1}\leq\frac{1}{2}\left\|\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{1},M)}_{\text{QDP}}(\sigma_{1})-\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{1},M)}_{\text{QDP}}(\rho_{1})\right\|_{1}+\frac{1}{2}\left\|\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{1},M)}_{\text{QDP}}-\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho_{1})}\right\|_{\operatorname{Tr}}$
$\leq\Xi_{1}+\frac{1}{2}\left\|\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{1},M)}_{\text{QDP}}-\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho_{1},M)}_{\text{QDP}}\right\|_{\operatorname{Tr}}+\frac{1}{2}\left\|\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho_{1},M)}_{\text{QDP}}-\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho_{1})}\right\|_{\operatorname{Tr}}=\mathcal{O}(M\Xi_{1}).$ (26)
@@ II.1 Locally accurate implementation and unfolding of the dynamic unitary using black box evolutions
In this subsection, we demonstrate how to locally accurately implement a recursion unitary, as in Eq. (19), given access to black box evolutions with respect to the instruction state. In particular, we consider the generalized recursion
$\hat{U}^{(\{f\},\rho)}=\hat{V}_{L}e^{if_{L}(\rho)}\hat{V}_{L-1}e^{if_{L-1}(\rho)}\cdots\hat{V}_{1}e^{if_{1}(\rho)}\hat{V}_{0}\ ,$ (27)
@@
$e^{-it\rho_{n}^{\otimes d}}=\hat{U}^{\otimes d}e^{-it\rho_{0}^{\otimes d}}\left(\hat{U}^{\dagger}\right)^{\otimes d}\ ,$ (28)
where $\hat{U}=\hat{U}^{(\{f\},\rho_{n-1})}\hat{U}^{(\{f\},\rho_{n-2})}\cdots\hat{U}^{(\{f\},\rho_{0})}$.
@@
In other words, a locally accurate implementation ^$\hat{\mathcal{E}}^{(\{\mathcal{N}\},\sigma_{k},M)}_{\text{BB}}(\sigma_{k})$, such that
$\frac{1}{2}\left\lVert\hat{\mathcal{E}}^{(\{\mathcal{N}\},\sigma_{k},M)}_{\text{BB}}(\sigma_{k})-\hat{\mathbf{E}}^{(\{\mathcal{N}\},\sigma_{k})}(\sigma_{k})\right\rVert_{1}\leq\epsilon\ ,$ (29)
@@
$P_{s_{j}}^{(j)}=\left(\prod_{i=1}^{s_{j}}A_{i}^{(j)}\rho\right)A_{0}^{(j)},$ (30)
@@
$\mathcal{N}_{s_{j}}^{(j)}\left(\bigotimes_{k=1}^{d}|{v_{k}}\rangle\!\langle{w_{k}}|_{k}\right)=A^{(j)}_{s_{j}}|{v_{s_{j}}}\rangle\!\langle{w_{s_{j}}}|_{s_{j}}A^{(j)}_{s_{j}-1}\cdots A^{(j)}_{1}|{v_{1}}\rangle\!\langle{w_{1}}|_{1}A^{(j)}_{0}.$ (31)
for any matrix element $\bigotimes_{k=1}^{d}|{v_{k}}\rangle\!\langle{w_{k}}|_{k}$, where each $v_{k},w_{k}\in\{1,\cdots,\dim(\rho)\}$. Then, it immediately gives $\mathcal{N}_{s_{j}}^{(j)}(\rho^{\otimes d})=P_{s_{j}}^{(j)}(\rho)$. The summation of such linear maps, $\mathcal{N}=\sum_{j=1}^{J}\mathcal{N}_{s_{j}}^{(j)}$ is also a linear map and yields the desired result $\mathcal{N}(\rho^{\otimes d})=f(\rho)$. The second part of the Lemma can be shown similarly. If $f(\mathmybb{1})\propto\mathmybb{1}$, then $\mathcal{N}(\mathmybb{1}^{\otimes d})=f(\mathmybb{1})\propto\mathmybb{1}$. If $f$ is Hermitian-preserving, we can make a decomposition where each $P_{s_{j}}^{(j)}$ and thus the construction Eq. (31) are also Hermitian-preserving. ∎
Combining Lemmas 3 and 4, we can implement $e^{-if(\rho)}$ with an arbitrarily small error by making queries to $e^{-i\rho^{\otimes d}t}$. Hence, the dynamic unitary Eq. (27) with $L$ such memory calls can also be locally accurately implemented with error $\epsilon$ by making $\mathcal{O}(1/\epsilon)$ queries to $e^{-i\rho^{\otimes d}t}$.
@@ II.2 Hermitian-preserving map exponentiation and its extension
^s$\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho)}_{s}(\sigma)=\operatorname{Tr}_{1}\left[e^{-i\hat{N}s}\left(\rho\otimes\sigma\right)e^{i\hat{N}s}\right]\ ,$ (32)
@@
It is useful to note that in general, the reference state $\rho$ and the target state $\sigma$ in Def. 5 need not be of the same dimension. The map $\hat{\mathcal{N}}$ is in general a map from $\mathcal{B}(\mathcal{H}^{\prime})$ to $\mathcal{B}(\mathcal{H})$ with $\dim(\mathcal{H})\neq\dim(\mathcal{H}^{\prime})$. In such cases, the reference state $\rho\in\mathcal{S}(\mathcal{H}^{\prime})$, while the output of the map $\mathcal{N}(\rho)\in\mathcal{B}(\mathcal{H})$ and the target state $\sigma\in\mathcal{S}(\mathcal{H})$. Likewise, the operator $\hat{N}\in\mathcal{B}(\mathcal{H}^{\prime}\otimes\mathcal{H})$, since the maximally entangled state used to define the Choi matrix $|{\Phi^{+}}\rangle_{12}\in\mathcal{H}^{\prime}\otimes\mathcal{H}^{\prime}$. The channel Eq. (32) well approximates the unitary channel
$\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho)}_{s}(\sigma):=e^{is\hat{\mathcal{N}}(\rho)}(\sigma)e^{-is\hat{\mathcal{N}}(\rho)}.$ (34)
@@
^s$\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho)}_{s}(\sigma)$ $=\sigma-is\operatorname{Tr}_{1}\left(\left[\hat{N},\rho\otimes\sigma\right]\right)+\mathcal{O}(s^{2})=\sigma-is\operatorname{Tr}_{1}\left(\left[\sum_{j,k}\left(|{k}\rangle\!\langle{j}|_{1}\otimes\hat{\mathcal{N}}\left(|{j}\rangle\!\langle{k}|\right)_{2}\right),\rho\otimes\sigma\right]\right)+\mathcal{O}(s^{2})\ .$ (35)
@@
^s$\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho)}_{s}(\sigma)$ $=\sigma-is\sum_{jk}\left[\operatorname{Tr}_{1}\left(|{k}\rangle\!\langle{j}|_{1}\rho\otimes\hat{\mathcal{N}}\left(|{j}\rangle\!\langle{k}|\right)_{2}\sigma\right)-\operatorname{Tr}_{1}\left(\rho|{k}\rangle\!\langle{j}|_{1}\otimes\sigma\hat{\mathcal{N}}\left(|{j}\rangle\!\langle{k}|\right)_{2}\right)\right]+\mathcal{O}(s^{2})$ (36)
@@
^s$\left\lVert\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho)}_{s}-\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho)}_{s}\right\rVert_{\diamond}\leq 8\|\hat{N}\|_{\infty}^{2}s^{2}\ ,$ (37)
whenever $\|\hat{N}\|_{\infty}s\in(0,0.8]$. Setting $s\mapsto s/M$ and repeating the channel $M$ times, we can better approximate $\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho)}_{s}$ and achieve
$\frac{1}{2}\left\lVert\left(\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho)}_{s/M}\right)^{M}-\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho)}_{s}\right\rVert_{\operatorname{Tr}}\leq\frac{1}{2}\left\lVert\left(\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho)}_{s/M}\right)^{M}-\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho)}_{s}\right\rVert_{\diamond}\leq 4\lVert\hat{N}\rVert_{\infty}^{2}\frac{s^{2}}{M}=\mathcal{O}\left(\lVert\hat{N}\rVert_{\infty}^{2}\frac{s^{2}}{M}\right)\ ,$ (38)
using the subadditivity (^$\|(\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho)}_{s/M})^{M}-\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho)}_{s}\|\leq M\|\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho)}_{s/M}-\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho)}_{s/M}\|$. Note that the $M$ applications of ^$\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho)}_{s/M}$ costs $M$ copies of $\rho$. Then Eq. (38) can be recast as the following statement: the Hermitian-preserving map exponentiation unitary $\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho)}_{s}$ can be implemented with an arbitrarily small error $\epsilon$ by consuming $\mathcal{O}(\lVert\hat{N}\rVert_{\infty}^{2}s^{2}/\epsilon)$ copies of $\rho$.
@@
^s$\hat{\mathcal{E}}^{(\rho)}_{s}(\sigma)=\operatorname{Tr}_{1}\left[e^{-i\hat{S}s}\left(\rho\otimes\sigma\right)e^{i\hat{S}s}\right]\ ,$ (39)

### v2

@@ II Locally accurate implementations
In Supplemental Materials, we focus on quantum recursions whose initial states are in general mixed. Then a memory-call unitary Eq. (5) becomes
$\sigma\mapsto\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho]}(\sigma)\coloneq e^{i\hat{\mathcal{N}}(\rho)}\sigma e^{-i\hat{\mathcal{N}}(\rho)}\ ,$ (S5)
with a memory state $\rho$, a working state $\sigma$, and a hermitian-preserving map $\hat{\mathcal{N}}$.
From this, we can define quantum recursions. Consider the simplest case with only one memory-call unitary of the form Eq. (S5) and trivial constant unitaries $V_{0}=V_{1}=\mathmybb{1}$ when written in the form Eq. (2). Starting from the root state $\rho_{0}$, the recursion is defined by $\rho_{n+1}=\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho_{n}]}(\rho_{n})$; the exact implementation would lead to a sequence of states $\rho_{0},\rho_{1},\ldots$ For both unfolding and QDP implementations, some memory-calls cannot be executed without error, and we need to consider the approximate implementations.
@@
$\frac{1}{2}\left\lVert\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{k},M]}_{\text{QDP}}({\sigma_{k}})-\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\sigma_{k}]}(\sigma_{k})\right\rVert_{1}=\frac{1}{2}\left\lVert\sigma_{k+1}-\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\sigma_{k}]}(\sigma_{k})\right\rVert_{1}\leq\epsilon.$ (S6)
@@
For clarity we assume the recursion step consists of a single memory-call implemented by $M$ memory-usage queries. We then have that the ideal implementation is $\rho_{1}=\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho_{0}]}(\rho_{0})=\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\sigma_{0}]}(\sigma_{0})$ and the practical implementation $\sigma_{1}=\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{0},M]}_{\text{QDP}}(\sigma_{0})$, such that
$\Xi_{1}$ :-$\coloneq\frac{1}{2}\|\sigma_{1}-\rho_{1}\|_{1}=\frac{1}{2}\left\|\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{0},M]}_{\text{QDP}}(\sigma_{0})-\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\sigma_{0}]}(\sigma_{0})\right\|_{1}=\mathcal{O}(1/M).:$ (S7)
@@
$\frac{1}{2}\left\|\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{0},M]}_{\text{QDP}}-\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\sigma_{0}]}\right\|_{\operatorname{Tr}}=\mathcal{O}(1/M)\ .$ (S8)
The QDP channel ^$\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{1},M]}_{\text{QDP}}$ consists of $M$ memory-usage queries. For each such memory-usage query defined by
^s$\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{1}]}_{s}(\omega)=\operatorname{Tr}_{1}\left[e^{-i\hat{N}s}\left(\sigma_{1}\otimes\omega\right)e^{i\hat{N}s}\right]\ ,$ (S9)
@@
^s$\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{1}]}_{s}(\omega)=\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho_{1}]}_{s}(\omega)+\operatorname{Tr}_{1}\left[e^{-i\hat{N}s}\left((\sigma_{1}-\rho_{1})\otimes\omega\right)e^{i\hat{N}s}\right]\ ,$ (S10)
@@
$\frac{1}{2}\left\|\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{1}]}_{s}-\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho_{1}]}_{s}\right\|_{\operatorname{Tr}}\leq\frac{1}{2}\|\sigma_{1}-\rho_{1}\|_{1}=\Xi_{1}\ .$ (S11)
Here we see how the preparation error $\Xi_{1}$ of $\sigma_{1}$ compared to $\rho_{1}$ influences the distance of the recursion step maps that define $\sigma_{2}=\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{1},M]}_{\text{QDP}}(\sigma_{1})$ and $\rho_{2}=\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho_{1},M]}_{\text{QDP}}(\rho_{1})$ so that using telescoping and the triangle inequality
$\frac{1}{2}\left\|\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{1},M]}_{\text{QDP}}-\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho_{1},M]}_{\text{QDP}}\right\|_{\operatorname{Tr}}\leq\frac{M}{2}\left\|\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{1}]}_{1/M}-\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho_{1}]}_{1/M}\right\|_{\operatorname{Tr}}=\mathcal{O}(M\Xi_{1})\ .$ (S12)
@@
$\Xi_{2}$ $=\frac{1}{2}\left\|\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{1},M]}_{\text{QDP}}(\sigma_{1})-\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho_{1}]}(\rho_{1})\right\|_{1}\leq\frac{1}{2}\left\|\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{1},M]}_{\text{QDP}}(\sigma_{1})-\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{1},M]}_{\text{QDP}}(\rho_{1})\right\|_{1}+\frac{1}{2}\left\|\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{1},M]}_{\text{QDP}}-\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho_{1}]}\right\|_{\operatorname{Tr}}$
$\leq\Xi_{1}+\frac{1}{2}\left\|\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{1},M]}_{\text{QDP}}-\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho_{1},M]}_{\text{QDP}}\right\|_{\operatorname{Tr}}+\frac{1}{2}\left\|\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho_{1},M]}_{\text{QDP}}-\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho_{1}]}\right\|_{\operatorname{Tr}}=\mathcal{O}(M\Xi_{1}).$ (S13)
@@ II.1 Locally accurate implementation and unfolding of the dynamic unitary using black box evolutions
In this subsection, we demonstrate how to locally accurately implement a recursion unitary, as in Eq. (S6), given access to black box evolutions with respect to the instruction state. In particular, we consider the generalized recursion
$\hat{U}^{[\{f\},\rho]}=\hat{V}_{L}e^{if_{L}(\rho)}\hat{V}_{L-1}e^{if_{L-1}(\rho)}\cdots\hat{V}_{1}e^{if_{1}(\rho)}\hat{V}_{0}\ ,$ (S14)
@@
$e^{-it\rho_{n}^{\otimes d}}=\hat{U}^{\otimes d}e^{-it\rho_{0}^{\otimes d}}\left(\hat{U}^{\dagger}\right)^{\otimes d}\ ,$ (S15)
where $\hat{U}=\hat{U}^{[\{f\},\rho_{n-1}]}\hat{U}^{[\{f\},\rho_{n-2}]}\cdots\hat{U}^{[\{f\},\rho_{0}]}$.
@@
In other words, a locally accurate implementation ^$\hat{\mathcal{E}}^{[\{\mathcal{N}\},\sigma_{k},M]}_{\text{BB}}(\sigma_{k})$, such that
$\frac{1}{2}\left\lVert\hat{\mathcal{E}}^{[\{\mathcal{N}\},\sigma_{k},M]}_{\text{BB}}(\sigma_{k})-\hat{\mathbf{E}}^{[\{\mathcal{N}\},\sigma_{k}]}(\sigma_{k})\right\rVert_{1}\leq\epsilon\ ,$ (S16)
@@
$P_{s_{j}}^{(j)}=\left(\prod_{i=1}^{s_{j}}A_{i}^{(j)}\rho\right)A_{0}^{(j)},$ (S17)
@@
$\mathcal{N}_{s_{j}}^{(j)}\left(\bigotimes_{k=1}^{d}|{v_{k}}\rangle\!\langle{w_{k}}|_{k}\right)=A^{(j)}_{s_{j}}|{v_{s_{j}}}\rangle\!\langle{w_{s_{j}}}|_{s_{j}}A^{(j)}_{s_{j}-1}\cdots A^{(j)}_{1}|{v_{1}}\rangle\!\langle{w_{1}}|_{1}A^{(j)}_{0}.$ (S18)
for any matrix element $\bigotimes_{k=1}^{d}|{v_{k}}\rangle\!\langle{w_{k}}|_{k}$, where each $v_{k},w_{k}\in\{1,\cdots,\dim(\rho)\}$. Then, it immediately gives $\mathcal{N}_{s_{j}}^{(j)}(\rho^{\otimes d})=P_{s_{j}}^{(j)}(\rho)$. The summation of such linear maps, $\mathcal{N}=\sum_{j=1}^{J}\mathcal{N}_{s_{j}}^{(j)}$ is also a linear map and yields the desired result $\mathcal{N}(\rho^{\otimes d})=f(\rho)$.
The second part of the Lemma can be shown similarly. If $f(\mathmybb{1})\propto\mathmybb{1}$, then $\mathcal{N}(\mathmybb{1}^{\otimes d})=f(\mathmybb{1})\propto\mathmybb{1}$. If $f$ is Hermitian-preserving, we can make a decomposition where each $P_{s_{j}}^{(j)}$ and thus the construction Eq. (S18) are also Hermitian-preserving. ∎
Combining Lemmas 3 and 4, we can implement $e^{-if(\rho)}$ with an arbitrarily small error by making queries to $e^{-i\rho^{\otimes d}t}$. Hence, the dynamic unitary Eq. (S14) with $L$ such memory calls can also be locally accurately implemented with error $\epsilon$ by making $\mathcal{O}(1/\epsilon)$ queries to $e^{-i\rho^{\otimes d}t}$.
@@ II.2 Hermitian-preserving map exponentiation and its extension
^s$\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho]}_{s}(\sigma)=\operatorname{Tr}_{1}\left[e^{-i\hat{N}s}\left(\rho\otimes\sigma\right)e^{i\hat{N}s}\right]\ ,$ (S19)
@@
It is useful to note that in general, the reference state $\rho$ and the target state $\sigma$ in Def. 5 need not be of the same dimension. The map $\hat{\mathcal{N}}$ is in general a map from $\mathcal{B}(\mathcal{H}^{\prime})$ to $\mathcal{B}(\mathcal{H})$ with $\dim(\mathcal{H})\neq\dim(\mathcal{H}^{\prime})$. In such cases, the reference state $\rho\in\mathcal{S}(\mathcal{H}^{\prime})$, while the output of the map $\mathcal{N}(\rho)\in\mathcal{B}(\mathcal{H})$ and the target state $\sigma\in\mathcal{S}(\mathcal{H})$. Likewise, the operator $\hat{N}\in\mathcal{B}(\mathcal{H}^{\prime}\otimes\mathcal{H})$, since the maximally entangled state used to define the Choi matrix $|{\Phi^{+}}\rangle_{12}\in\mathcal{H}^{\prime}\otimes\mathcal{H}^{\prime}$. The channel Eq. (S19) well approximates the unitary channel
$\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho]}_{s}(\sigma):=e^{is\hat{\mathcal{N}}(\rho)}(\sigma)e^{-is\hat{\mathcal{N}}(\rho)}.$ (S21)
@@
^s$\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho]}_{s}(\sigma)$ $=\sigma-is\operatorname{Tr}_{1}\left(\left[\hat{N},\rho\otimes\sigma\right]\right)+\mathcal{O}(s^{2})=\sigma-is\operatorname{Tr}_{1}\left(\left[\sum_{j,k}\left(|{k}\rangle\!\langle{j}|_{1}\otimes\hat{\mathcal{N}}\left(|{j}\rangle\!\langle{k}|\right)_{2}\right),\rho\otimes\sigma\right]\right)+\mathcal{O}(s^{2})\ .$ (S22)
@@
^s$\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho]}_{s}(\sigma)$ $=\sigma-is\sum_{jk}\left[\operatorname{Tr}_{1}\left(|{k}\rangle\!\langle{j}|_{1}\rho\otimes\hat{\mathcal{N}}\left(|{j}\rangle\!\langle{k}|\right)_{2}\sigma\right)-\operatorname{Tr}_{1}\left(\rho|{k}\rangle\!\langle{j}|_{1}\otimes\sigma\hat{\mathcal{N}}\left(|{j}\rangle\!\langle{k}|\right)_{2}\right)\right]+\mathcal{O}(s^{2})$ (S23)
@@
^s$\left\lVert\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho]}_{s}-\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho]}_{s}\right\rVert_{\diamond}\leq 8\|\hat{N}\|_{\infty}^{2}s^{2}\ ,$ (S24)
whenever $\|\hat{N}\|_{\infty}s\in(0,0.8]$. Setting $s\mapsto s/M$ and repeating the channel $M$ times, we can better approximate $\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho]}_{s}$ and achieve
$\frac{1}{2}\left\lVert\left(\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho]}_{s/M}\right)^{M}-\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho]}_{s}\right\rVert_{\operatorname{Tr}}\leq\frac{1}{2}\left\lVert\left(\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho]}_{s/M}\right)^{M}-\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho]}_{s}\right\rVert_{\diamond}\leq 4\lVert\hat{N}\rVert_{\infty}^{2}\frac{s^{2}}{M}=\mathcal{O}\left(\lVert\hat{N}\rVert_{\infty}^{2}\frac{s^{2}}{M}\right)\ ,$ (S25)
using the subadditivity (^$\|(\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho]}_{s/M})^{M}-\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho]}_{s}\|\leq M\|\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho]}_{s/M}-\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho]}_{s/M}\|$. Note that the $M$ applications of ^$\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\rho]}_{s/M}$ costs $M$ copies of $\rho$. Then Eq. (S25) can be recast as the following statement: the Hermitian-preserving map exponentiation unitary $\hat{\mathbf{E}}^{[\hat{\mathcal{N}},\rho]}_{s}$ can be implemented with an arbitrarily small error $\epsilon$ by consuming $\mathcal{O}(\lVert\hat{N}\rVert_{\infty}^{2}s^{2}/\epsilon)$ copies of $\rho$.
@@
^s$\hat{\mathcal{E}}^{[\rho]}_{s}(\sigma)=\operatorname{Tr}_{1}\left[e^{-i\hat{S}s}\left(\rho\otimes\sigma\right)e^{i\hat{S}s}\right]\ ,$ (S26)